The sum of ages of father nd son is 56.Four years ago the father was 3times as old as son. Find thier present age.
Answer by jorel555(1290)

Let father and son be f and s. Then:
f+s=56
f-4=3(s-4). So:
f-4=3s-12
f=3s-8
3s-8+s=56
4s=64
s=16
f=40
The father is now 40; the son is now 16. ☺☺☺☺